Determine the pH of an aqueous buffer solution (250 mL) comprised of dissolving 5.2812 g of Tris (B) (FW = 121.14 g/mol) and 10.2500 g of Tris-HCl (BH+) (FW = 157.60 g/mol) if pKa (BH+) = 8.07?

(a) 8.24 (b) 7.90 (c) 8.54 (d) 7.82 (e) 8.01
Answer #1

The pH can be calculated using the Henderson-Hasselbach equation:

[B] pH = pka + log BH+1)

We need to calculate the concentrations of the species, which can be done using:

M =mmolar V(L)

For B:

5.28129 1B1 - 121.14g/mol 0.2501 = 0.174.M

For BH+:
BH+1 -157.60g/mol 10.2500g -= 0.260M 0.250L

And finally, we can calculate:

0.174M pH = 8.07 + log 7.90 0.260M